OpenStudy (anonymous):

The 1st electron in the H2 atom is ejected from the atom after absorbing a total amt of Energy = 20.4eV . Calculate its de Broglie wavelenght in A^0

OpenStudy (anonymous):

@UnkleRhaukus @mukushla @hartnn

OpenStudy (anonymous):

\[E=hc/\lambda\] Does that help?

OpenStudy (anonymous):

i tried this,,,but did nt get...though

OpenStudy (anonymous):

E = 20.4 - 13.6 = 6.8eV

OpenStudy (anonymous):

*re-check calculation* Because as far as I know, De-broglie's eqn has been well tested ~~

OpenStudy (unklerhaukus):

\[c=\nu\lambda\qquad \qquad\nu=\frac c\lambda\] \[E=h\nu\qquad\qquad E=\frac{hc}\lambda\] \[\lambda=\frac{hc}{E}\] \[E = 20.4[\text{eV}]\]\[h=4.14×10^{−15}\left[\text{eV}\cdot\text {s}\right]\]\[c=3.00\times10^8\left[\frac{\text m}{\text s}\right]\] \[\lambda=\frac{4.14×10^{−15}\left[\text{eV}\cdot\text {s}\right]\times3.00\times10^8\left[\frac{\text m}{\text s}\right]}{20.4[\text{eV}]}\]

OpenStudy (anonymous):

^no, that doesn't take into account the binding energy required to liberate the electron.

OpenStudy (unklerhaukus):

\[609Å\] what have i forgotten Jemurray3?

OpenStudy (anonymous):

It absorbs 20.4 eV, but some of that goes into liberating it from the molecule.

OpenStudy (unklerhaukus):

So \(E=20.4[\text {eV}]-\text {binding energy of electron}\)

OpenStudy (unklerhaukus):

is that 13.7 eV?

OpenStudy (anonymous):

it is E = 20.4 - 13.6 = 6.8eV

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

To clarify, when you say H2 atom, you mean molecular hydrogen, right? Like, two hydrogen atoms bound together?